# CurbNet: Curb Detection Framework Based on LiDAR Point Cloud Segmentation

### v1

@@ III-D Loss Group
In the context of curb detection for autonomous driving, we propose a novel combination of loss functions to adapt to the class imbalance in training data. In real-world scenarios, the point cloud data for curbs comprises only a small fraction compared to other categories such as roads and buildings. Using uniform loss weights can lead to training imbalances, adversely affecting the recognition performance for the minority class, i.e., curbs. To address this, we introduce a combination of Adaptive Cross-Entropy (CE) Loss and Lovász-Softmax Loss.
1) Adaptive CE Loss
This paper improves upon the standard Cross-Entropy (CE) Loss by adjusting weights based on the distribution of point cloud data across different classes, enabling the model to focus more on learning from scarce sample data. The standard CE Loss is defined as:
$\mathcal{L}_{CE}=-y_{1}*\log\left(p_{1}\right)-y_{2}*\log\left(p_{2}\right)$ (1)
where $y_{i}$ represents the true label and $p_{i}$ represents the predicted probability for class $i$.
To address class imbalance, we modify the loss weights for different classes using the $\alpha$ parameter, akin to the weighted CE Loss. In this context, $\alpha$ is computed based on the inverse ratio of point cloud quantities of each class, followed by normalization. The formulation of the Weighted CE Loss ($\mathcal{L}_{WCE}$) is:
$\mathcal{L}_{WCE}=-\alpha_{1}*y_{1}*\log\left(p_{1}\right)-\alpha_{2}*y_{2}\log\left(p_{2}\right)$ (2)
where $\alpha_{i}$ is the class-specific weight calculated as:
$\alpha_{i}=\frac{1}{\log(\delta+\frac{N}{N_{i}})}$ (3)
Here, $N$ is the total number of points in the point cloud, $N_{i}$ is the number of points in class $i$, and $\delta$ is a small constant to prevent division by zero.
Further drawing inspiration from Focal Loss, we introduce a modulation factor $(1-P_{t})^{\gamma}$ into the cross-entropy loss, effectively reducing the weights for easily classified samples and increasing them for the more challenging ones, thereby directing the training focus towards difficult-to-classify samples. The Adaptive CE Loss ($\mathcal{L}_{ACE}$) is therefore formulated as:
$\begin{split}\mathcal{L}_{ACE}=&-\alpha_{1}*(1-p_{1})^{\gamma}*y_{1}*\log(p_{1})\\ &-\alpha_{2}*(1-p_{2})^{\gamma}*y_{2}*\log(p_{2})\end{split}$ (4)
The focusing parameter $\gamma$, which modulates the emphasis on difficult samples, is set to 2 in this study.
@@
Lovász Loss is particularly effective in handling imbalanced datasets and excels in addressing sparse boundary issues[45]. Compared to traditional cross-entropy loss, it demonstrates superior performance in terms of Intersection over Union (IoU) scores, commonly used for evaluating image segmentation results due to their quality perception and scale invariance. In Lovász Loss, for a given true label vector $\boldsymbol{y}^{*}$ and a predicted label vector $\widetilde{\boldsymbol{y}}$, the IoU index for class c is defined as:
@@
The Lovász-Softmax loss extends this concept by applying the Lovász extension to the softmax probabilities of a model’s output. It optimizes a convex surrogate of the IoU score, which is more suitable for gradient-based optimization. Specifically, the Lovász-Softmax loss $\mathcal{L}_{IoU}$ for a set of classes $C$ is defined as:
$\mathcal{L}_{(}\boldsymbol{y}^{*},\widetilde{\boldsymbol{y}})=\sum_{c\in C}\Delta_{\text{IoU}_{c}}\left(\boldsymbol{y}^{*},\widetilde{\boldsymbol{y}}\right)$ (7)
The computation involves ordering the pixels by their error margin and computing a weighted sum of the individual errors, thus directly targeting the errors that most impact the IoU score.
In practice, the Lovász-Softmax loss is particularly beneficial for segmenting objects with irregular or sparse boundaries, as it focuses on the segments of the predictions that have the largest impact on the overall IoU score, rather than treating all errors equally.
@@ III-E Multi-Cluster and Curve Fitting
where $\operatorname{dist}(P,Q)$ represents the distance between points P and Q. A point P is considered a core point if its $\varepsilon$-neighborhood contains at least minPts points, i.e., $\left|N_{\varepsilon}(P)\right|\geq\operatorname{minPts}$.
@@
During the polynomial curve fitting of curbs, the aim is to eliminate noise points not belonging to the curb. Assuming the curve equation is $f(x)$, for any point $P(x,y)$ in the point cloud, we calculate the perpendicular distance $d$ from the point to the curve:
@@
The complete operation process is shown in Algorithm. 1. Through this approach, combining the DBSCAN algorithm with polynomial curve fitting effectively identifies and extracts accurate curb lines from LiDAR point cloud data, while eliminating noise points, thus improving the overall segmentation accuracy.
@@
3:Step 1: Apply DBSCAN to point cloud data
@@
5:     Compute $N_{\varepsilon}(P)$
@@
11:Step 3: Fit polynomial curves to each curb segment
@@
16:         if $d(P,f)>\delta$ then
17:              Remove $P$ as noise
18:         end if
19:     end for
@@
21:Step 4: Output refined curb line segments
@@ IV-A Experiment Setup
Our model training was conducted in an Ubuntu 20.04 environment, utilizing an Intel(R) Xeon(R) Gold 5318S CPU @ 2.10GHz and an NVIDIA RTX 3090 GPU. We employed the PyTorch framework for model training, using the Adam optimizer. The training parameters were set with a batch size of 8 and a total of 100 epochs, at a learning rate of 0.001.
@@
The performance of our curb detection method was rigorously evaluated using standard metrics. These include Precision, Recall, and F-1 score, which are quintessential for quantifying the accuracy and reliability of classification models. Precision, defined as $\frac{TP}{TP+FP}$, measures the proportion of correctly predicted positive observations to the total predicted positives. Recall, calculated as $\frac{TP}{TP+FN}$, assesses the proportion of actual positives that were correctly identified. The F-1 score, given by $2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$, harmonizes the balance between Precision and Recall, providing a single measure of efficacy. Here, TP (True Positives) represents the number of correct positive predictions, FP (False Positives) denotes the count of negative instances incorrectly classified as positive, TN (True Negatives) refers to the count of correct negative predictions, and FN (False Negatives) signifies the instances where positive cases were wrongly predicted as negative. These metrics collectively offer a comprehensive view of our model’s performance, crucial for its validation in high-level applications.
@@ IV-B Quantitative Results of Curb Detection
In the NRS dataset experiments (refer to Table. II), this study compared classic segmentation algorithms such as PointPillars, U-Net, Swin-Transformer, and CSWin-Transformer, as well as the state-of-the-art curb detection model, LCDeT. Thanks to the specially designed MSCA module for Curb 3D scenarios, our proposed CurbNet achieved the highest detection performance on the NRS dataset, with Precision, Recall, and F-1 scores reaching 0.8280, 0.8341, and 0.8309, respectively.
TABLE II: COMPARISON OF RESULTS IN NRS-DATASET[25].
@@
CurbNet (ours) 0.8280 0.8341 0.8309
CurbNet-post (ours) 0.8420 0.8497 0.8458
The experiments in the 3D-Curb dataset not only compared classic and advanced deep learning model algorithms but also included three traditional methods of manual feature extraction, as shown in Table. III. Owing to the robust automatic feature extraction capabilities, deep learning methods significantly outperformed in curb detection, surpassing the other methods by over 10 points in Precision, Recall, and F-1 score. Among the deep learning methods, our CurbNet surpassed the best-performing supervised learning model on the SemantickITTI dataset, Cylinder3D, by more than 2 points in Precision. This demonstrates the exceptional performance of CurbNet in the field of curb point cloud segmentation.
TABLE III: COMPARISON OF RESULTS IN 3D-Curb DATASET.
@@
CurbNet (ours) 0.8293 0.8569 0.8429
CurbNet-post (ours) 0.8744 0.8648 0.8696
@@
In our study, we conducted Tolerance performance tests on the 3D-Curb dataset, setting four Tolerances ranging from just 0.05m to 0.2m, as shown in Table. IV. Tests were carried out on three models: 3D U-Net, Cylinder 3D, and CurbNet (ours). With the increase in error Tolerance, performance metrics improved significantly. At 0.05m Tolerance, Precision improved by an average of 4 points; at 0.1m, by 9 points; at 0.15m, by 12 points; and at 0.2m, by 13 points. As Tolerance increased, performance gains gradually reached saturation. Notably, our CurbNet exceeded 0.95 in the average values of Precision, Recall, and F-1 score at just 0.15m Tolerance. This represents the optimal performance achieved in curb detection based on point cloud segmentation.
TABLE IV: Comparison of different tolerances in 3D-Curb Dataset.
@@
CurbNet (ours) 0.8665 0.9160 0.9460 0.9589 0.9031 0.9433 0.9664 0.9761 0.8844 0.9294 0.9561 0.9674
CurbNet-post (ours) 0.9139 0.9532 0.9682 0.9740 0.8974 0.9301 0.9439 0.9501 0.9056 0.9415 0.9559 0.9619
@@ IV-C Visualization Results of Curb Detection
We conducted a visual analysis of the test results obtained using our CurbNet model, comparing them with the ground truth to further validate the model’s detection performance. Fig. 5 illustrates the visualization results in scenarios without obstructions, showcasing the curb detection results in five common road scenes: straight road, curved road, right-angle intersection, curved intersection, and cross intersection. The images clearly demonstrate that the CurbNet model successfully identified all curb lines present in the ground truth, even in complex intersection scenarios. Notably, in the curved intersection scenario, the results identified by CurbNet were even more precise than the ground truth annotations, showcasing our model’s exceptional feature extraction capabilities and generalizability.
@@
Figure 6: Curb detection result in 3D-Curb dataset under occlusion. We compared the curb detection results at five classic intersections with occlusion. Even if there is occlusion, it does not affect the curb detection in other areas at all.
Furthermore, we conducted a visual analysis of the results on the NRS dataset, as shown in Fig. 7. This analysis primarily showcases curb detection in five common road scenarios and four special intersection types. Through a comparative visualization of the detection results and ground truth, our proposed CurbNet accurately identified the respective curb features.
@@ IV-E Ablation Study
As shown in Table. V, this study conducted comparative ablation experiments focusing on the main loss functions and crucial module designs of our model. We first conducted individual experiments on the employed $\mathcal{L}_{CE}$ (Cross-Entropy), $\mathcal{L}_{ACE}$ (Adaptive Cross-Entropy), and $\mathcal{L}_{IoU}$ (Intersection over Union) losses. Subsequently, we tested combinations of $\mathcal{L}_{CE}$+$\mathcal{L}_{IoU}$ loss and $\mathcal{L}_{ACE}$+$\mathcal{L}_{IoU}$ loss. In the experimental results, due to the design of the $\mathcal{L}_{ACE}$ loss addressing the imbalance in the number of curb point clouds compared to other categories, its disproportionate weight settings caused the model to overly focus on recall during training. However, the interaction with the $\mathcal{L}_{IoU}$ loss led to a more balanced overall performance, achieving optimal detection capabilities. The $\mathcal{L}_{ACE}$+$\mathcal{L}_{IoU}$ loss combination outperformed the $\mathcal{L}_{CE}$+$\mathcal{L}_{IoU}$ loss group by 1.5 points in Precision and 2 points in Recall.
Finally, we compared the model’s performance with and without the MSCA module. Under the $\mathcal{L}_{ACE}$+$\mathcal{L}_{IoU}$ loss setting, including the MSCA module significantly enhanced the corresponding performance metrics, particularly increasing Recall by 2 points and F-1 score by 1 point.
@@
$\mathcal{L}_{CE}$ $\mathcal{L}_{ACE}$ $\mathcal{L}_{IoU}$ MSCA Precision Recall F-1 score

### v2

@@ III-D Loss Group
In real-world scenarios, the point cloud data for curbs comprises only a small fraction compared to other categories such as roads and buildings. Using uniform loss weights can lead to training imbalances, adversely affecting the recognition performance for the minority class, i.e., curbs. To address this, we propose a novel combination of Adaptive Cross-Entropy (ACE) Loss and Lovász-Softmax Loss.
1) Adaptive Cross Entropy (ACE) Loss
Due to the imbalance between the number of curb point clouds and the number of other categories such as roads and buildings. Standard loss functions like Cross Entropy (CE)[44] do not adequately address this imbalance, leading to suboptimal performance in recognizing minority classes. The standard CE loss is defined as:
$\mathcal{L}_{CE}\left(p_{\mathrm{t}}\right)=-\log\left(p_{\mathrm{t}}\right)$ (1)
where $p_{t}$ represents the predicted probability of the true class.
Given the disproportionate representation of classes in the point cloud data, we draw inspiration from the Focal Loss[45] to reallocate the loss contribution of easy and hard samples, significantly reducing the influence of the majority background samples:
$\mathcal{L}_{FL}\left(p_{\mathrm{t}}\right)=-\alpha_{\mathrm{t}}\left(1-p_{\mathrm{t}}\right)^{\gamma}\log\left(p_{\mathrm{t}}\right)$ (2)
The modulation factor $(1-p_{t})^{\gamma}$ in Focal Loss is crucial as it down-weights the loss for well-classified examples and focuses learning on hard examples. However, Focal Loss treats all classes equally with the same modulation factor, which does not address the imbalance among foreground classes.
Adaptive Class-wise Focusing Factor. To tackle both the foreground-background imbalance and the inter-foreground class imbalance, we introduce an adaptive, class-wise focusing factor $\gamma^{i}$ that adjusts according to the imbalance degree of each class $i$. The adaptive focusing factor $\gamma^{i}$ is defined as:
$\gamma^{i}$ $=\gamma_{a}+\gamma_{b}^{i}$ (3)
$=\gamma_{a}+s\left(1-\eta^{i}\right)$
Here, $\gamma^{i}$ is decomposed into a class-agnostic parameter $\gamma_{a}$ and a class-specific parameter $\gamma_{b}^{i}$. The parameter $\gamma_{a}$ represents the basic focusing factor under balanced data scenarios, while $\gamma_{b}^{i}\geq 0$ is a variable parameter related to the imbalance degree of class $i$. The term $\eta^{i}=\frac{N}{N_{i}}$ where $N$ is the total number of points in the point cloud, and $N_{i}$ is the number of points in class $i$. The value of $g^{i}$ is constrained to the range $[0,1]$, and $1-\eta^{i}$ inversely reflects the weight for low-frequency classes. The hyperparameter $s$ is a scaling factor that determines the upper limit of $\gamma^{i}$.
Dynamic Weight Factor. While the adaptive focusing factor $\gamma^{i}$ ensures more loss contribution from rare samples, it does not fully resolve the class imbalance problem. Therefore, we introduce a dynamic weighting factor $\omega^{i}$ to provide higher weights for rare classes:
$\omega^{i}=\frac{1}{\log(\delta+\eta^{i})}$ (4)
where $\delta$ is a small constant to prevent division by zero.
Combining these components, the final ACE Loss is expressed as:
$\mathcal{L}_{ACE}\left(p_{\mathrm{t}}\right)$ $=-\alpha_{\mathrm{t}}\omega^{i}\left(1-p_{\mathrm{t}}\right)^{\gamma^{i}}\log\left(p_{\mathrm{t}}\right)$ (5)
$=-\sum_{i=1}^{C}\alpha_{\mathrm{t}}\frac{1}{\log(\delta+\eta^{i})}\left(1-p_{\mathrm{t}}\right)^{\gamma_{a}+\gamma_{b}^{i}}\log\left(p_{\mathrm{t}}\right)$
The ACE Loss effectively prioritizes the learning of rare class samples by dynamically adjusting both the focusing factor and the class weights based on the distribution of point cloud data, thereby addressing the critical issue of class imbalance in curb detection.
@@
Lovász Loss is particularly effective in handling imbalanced datasets and excels in addressing sparse boundary issues[46]. Compared to traditional cross-entropy loss, it demonstrates superior performance in terms of Intersection over Union (IoU) scores. For a given true label vector $\boldsymbol{y}^{*}$ and a predicted label vector $\widetilde{\boldsymbol{y}}$, the IoU index for class c is defined as:
@@
The Lovász-Softmax loss extends this concept by applying the Lovász extension to the softmax probabilities of a model’s output. It optimizes a convex surrogate of the IoU score, which is more suitable for gradient-based optimization. Specifically, the loss $\mathcal{L}_{IoU}$ for a set of classes $C$ is defined as:
$\mathcal{L}_{IoU}(\boldsymbol{y}^{*},\widetilde{\boldsymbol{y}})=\sum_{c\in C}\Delta_{\text{IoU}_{c}}\left(\boldsymbol{y}^{*},\widetilde{\boldsymbol{y}}\right)$ (8)
The computation involves ordering the pixels by error margin and computing a weighted sum of the individual errors, thus directly targeting the errors that most impact the IoU score.
@@ III-E Multi-Cluster and Curve Fitting
where $\operatorname{dist}(P,Q)$ represents the distance between points P and Q. P is considered a core point if its $\varepsilon$-neighborhood contains at least minPts points, i.e., $\left|N_{\varepsilon}(P)\right|\geq\operatorname{minPts}$.
@@
During the polynomial curve fitting, the aim is to eliminate noise points not belonging to the curb. Assuming the curve equation is $f(x)$, for any point $P(x,y)$ in the point cloud, we calculate the perpendicular distance $d$ to the curve:
@@
The complete operation process is shown in Algorithm. 1. In order to improve calculation efficiency, we use parallel computing and KDTree’s efficient point cloud search scheme. Through this approach, combining the DBSCAN algorithm with polynomial curve fitting effectively identifies and extracts accurate curb lines, while eliminating noise points, thus improving the overall detection accuracy.
@@
3:Step 1: Apply DBSCAN to point cloud data with parallel processing
@@
5:     Compute $N_{\varepsilon}(P)$ in parallel computation
@@
11:Step 3: Fit polynomial curves in batch processing
@@
16:         Build KDTree for efficient nearest neighbor search
17:         if $d(P,f)>\delta$ then
18:              Remove $P$ as noise
19:         end if
@@
21:end for
22:Step 4: Output refined curb line segments
@@ IV-A Experiment Setup
Our model training was conducted in an Ubuntu 20.04 environment, utilizing an Intel(R) Xeon(R) Gold 5318S CPU @ 2.10GHz and an NVIDIA RTX 3090 GPU. We employed the PyTorch framework for model training and set training parameters with a batch size of 6, a total of 100 epochs, and a learning rate of 0.001.
@@
The performance of our curb detection method was rigorously evaluated using standard metrics. These include Precision, Recall, and F-1 score, which are quintessential for quantifying the accuracy and reliability of classification models. Precision, defined as $\frac{TP}{TP+FP}$, measures the proportion of correctly predicted positive observations to the total predicted positives. Recall, calculated as $\frac{TP}{TP+FN}$, assesses the proportion of actual positives that were correctly identified. The F-1 score, given by $2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$, harmonizes the balance between Precision and Recall, providing a single measure of efficacy. Here, TP (True Positives) represents the number of correct positive predictions, FP (False Positives) denotes the count of negative instances incorrectly classified as positive, TN (True Negatives) refers to the count of correct negative predictions, and FN (False Negatives) signifies the instances where positive cases were wrongly predicted as negative. These metrics collectively offer a comprehensive view of our model’s performance, which is crucial for the validation.
@@ IV-B Quantitative Results of Curb Detection
In the NRS dataset experiments (refer to Table. II), this study compared classic segmentation algorithms such as PointPillars, U-Net, Swin-Transformer, and CSWin-Transformer, as well as the state-of-the-art curb detection model, LCDeT. Leveraging the specially designed MSCA module for 3D curb scenarios, CurbNet achieved the highest detection performance on the NRS dataset. With the auxiliary training of relevant labels, CurbNet attained Precision, Recall, and F-1 scores of 0.8281, 0.8329, and 0.8308, respectively. Among them, auxiliary training helps improve Precision by 0.5 points, and post-processing helps improve Precision by 1.4 points.
TABLE II: COMPARISON OF RESULTS IN NRS-DATASET[25]. (w/o) and (w/) represent WITHOUT or WITH the auxiliary training of road and sidewalk labels respectively.
@@
CurbNet (w/o) 0.8225 0.8264 0.8234
CurbNet (w) 0.8281 0.8329 0.8308
CurbNet-post 0.8420 0.8496 0.8457
The experiments in the 3D-Curb dataset not only compared classic and advanced deep learning model algorithms but also included three traditional methods of manual feature extraction, as shown in Table. III. Owing to the robust automatic feature extraction capabilities, deep learning methods significantly outperformed in curb detection, surpassing the other methods by over 10 points in Precision, Recall, and F-1 score. Among the deep learning methods, CurbNet surpassed the best-performing supervised learning model on the SemanticKITTI dataset, Cylinder3D, by 2.5 points in Precision, and exceeded the knowledge distillation model PVKD by 1.5 points. Among them, Precision is improved by more than 1 point through auxiliary training, and by 4.5 points through post-processing.
TABLE III: COMPARISON OF RESULTS IN 3D-Curb DATASET. (w/o) and (w/) represent WITHOUT or WITH the auxiliary training of road and sidewalk labels respectively.
@@
PVKD [52] 0.8125 0.8089 0.8025
CurbNet (w/o) 0.8178 0.8395 0.8325
CurbNet (w/) 0.8292 0.8567 0.8427
CurbNet-post 0.8743 0.8647 0.8695
@@
In our study, we conducted Tolerance performance tests on the 3D-Curb dataset, setting four Tolerances ranging from just 0.05m to 0.2m, as shown in Table. IV. Tests were carried out on four models: 3D U-Net, Cylinder 3D, PVKD, and CurbNet (ours). With the increase in error Tolerance, performance metrics improved significantly. At 0.05m Tolerance, Precision improved by an average of 4 points; at 0.1m, by 9 points; at 0.15m, by 12 points; and at 0.2m, by 13 points. As Tolerance increased, performance gains gradually reached saturation. Notably, our CurbNet exceeded 0.95 in the average values of Precision, Recall, and F-1 score at just 0.15m Tolerance. This represents the optimal performance achieved in curb detection based on point cloud segmentation.
TABLE IV: Comparison of different tolerances in 3D-Curb Dataset. (w/o) and (w/) represent WITHOUT or WITH the auxiliary training of road and sidewalk labels respectively.
@@
PVKD [52] 0.8422 0.8898 0.9264 0.9398 0.8595 0.9021 0.9305 0.9417 0.8419 0.8902 0.9188 0.9343
CurbNet (w/o) 0.8588 0.9091 0.9397 0.9503 0.8934 0.9355 0.9559 0.9689 0.8793 0.9211 0.9476 0.9597
CurbNet (w/) 0.8667 0.9158 0.9461 0.9587 0.9029 0.9434 0.9663 0.9760 0.8845 0.9295 0.9562 0.9673
CurbNet-post 0.9139 0.9530 0.9679 0.9738 0.8976 0.9303 0.9441 0.9501 0.9056 0.9413 0.9558 0.9617
@@ IV-C Visualization Results of Curb Detection
We conducted a visual analysis of the test results obtained using our CurbNet model, comparing them with the ground truth to further validate the model’s detection performance. Fig. 6 illustrates the visualization results in scenarios without obstructions, showcasing the curb detection results in five common road scenes: straight road, curved road, right-angle intersection, curved intersection, and cross intersection. The images clearly demonstrate that the CurbNet model successfully identified all curb lines present in the ground truth, even in complex intersection scenarios. Notably, in the curved intersection scenario, the results identified by CurbNet were even more precise than the ground truth annotations, showcasing our model’s exceptional feature extraction capabilities and scene generalizability.
@@
Figure 7: Curb detection result in 3D-Curb dataset under occlusion. We compared the curb detection results at five classic intersections with occlusion. Even under occlusion, it does not affect curb detection in other areas.
Furthermore, we conducted a visual analysis of the results on the NRS dataset, as shown in Fig. 8. This analysis primarily showcases curb detection in five common road scenarios and four special intersection types. Through a comparative visualization of the detection results and ground truth, our method accurately identified the respective curb features.
@@ IV-E Ablation Study
As shown in Table. V, this study conducted comparative ablation experiments focusing on the main loss functions and crucial module designs of our model. We first conducted individual experiments on the employed $\mathcal{L}_{CE}$ (Cross-Entropy Loss), $\mathcal{L}_{FL}$ (Focal Loss), $\mathcal{L}_{ACE}$ (Adaptive Cross-Entropy Loss), and $\mathcal{L}_{IoU}$ (Intersection over Union Loss). Subsequently, we tested combinations of $\mathcal{L}_{CE}$+$\mathcal{L}_{IoU}$ loss, $\mathcal{L}_{FL}$+$\mathcal{L}_{IoU}$ loss and $\mathcal{L}_{ACE}$+$\mathcal{L}_{IoU}$ loss. In the experimental results, due to the design of the $\mathcal{L}_{ACE}$ loss addressing the imbalance in the number of curb point clouds compared to other categories, its disproportionate weight settings caused the model to overly focus on recall during training. However, the interaction with the $\mathcal{L}_{IoU}$ loss led to a more balanced overall performance, achieving optimal detection capabilities. The combination of $\mathcal{L}_{ACE}$+$\mathcal{L}_{IoU}$ loss outperformed the $\mathcal{L}_{CE}$+$\mathcal{L}_{IoU}$ loss group by 1.3 points in Precision and 2.2 points in Recall. It also surpassed the $\mathcal{L}_{FL}$+$\mathcal{L}_{IoU}$ loss group with improvements of 0.7 points in Precision and 2 points in Recall.
Finally, we compared the model’s performance with and without the MSCA module. Under the $\mathcal{L}_{ACE}$+$\mathcal{L}_{IoU}$ loss setting, including the MSCA module significantly enhanced the performance metrics, particularly increasing Recall by 2 points and the F-1 score by 1 point.
@@
$\mathcal{L}_{CE}$ $\mathcal{L}_{FL}$ $\mathcal{L}_{ACE}$ $\mathcal{L}_{IoU}$ MSCA Precision Recall F-1